INAF Pisces observing, 2011 November 29-30

Observer: C. Arcidiacono, G. Brusa, K. Boutsia, M. Fumana
Telescope Operator: G. Bec
Support Astronomer: -
Instrument: Pisces

Summary

Thick clouds on the whole sky late in the afternoon so we were not able to do flats. 15min after sunset the sky is clearing and at 01:00UT the sky is clear with cirrus in the horizon.
After 09:00UT (2AM MST) we have not been able to observe first because of an error with the guide probe (2hr lost) and then the AO was giving problems.
1hr lost to bad seeing that was varying between 2''-3''.
In total we lost 7hr for a variety of reasons. For more technical details see log in the AO twiki by G. Brusa at Log_20111130 .
In the remaining 5 hours we completed (90%) two projects:IC10 and NGC1560.

Data Summary

Project Objects Filter ExpTime Notes
Bono IC10 Ks , J 40min (+15min sky), 16min (+5min sky)  
Greggio NGC1560 Ks , J 42min, 38min  

Details

01:00 Slew for pointing correction and collimation

01:10 preset IC10
  • error: camera lense temperature is too low

01:15 slew to zenith because we need to switch off the treehouse

01:30 preset IC10 - again
  • there was an error in the script - first autocollimate and then preset
  • we collimated and then when the preset arrived found the mirror out of focus again (why??? - it shouldn't) and the runao failed
  • did the shell change when we sent the preset ??

01:45 repeat preset IC10
  • correct script this time - collimated after 5 iterations - runao
  • AO has sent a big correction to the mirror and the shell ripped
  • the same problem as yesterday - restart WFS Arbitrator gui - rerotator ON
  • AOS lost connection to Arbitrator - reboot AOS
  • the AOS has problems connecting to the telescope

01:57 repeat preset IC10
  • when the preset does not find the shell set it should: either set the shell or fail - here it just went on, while we were manually setting the shell
  • this should be fixed in software level.
  • dimm seeing 1.2''
  • it was not able to stop AO - commands kept coming but the mirror was not moving during gain optimization
  • cancel preset manually with AO stop

02:06 repeat preset IC10
  • during autogain - it started giving skip frames and then it failed

02:14 repeat preset IC10
  • this time the preset was successful

02:22 IC10 dither pattern in Ks
  • in today's images the background is lower than yesterday (700counts vs 800counts) and the sources have double the flux
  • we will try to repeat the whole program
  • dimm seeing 1.0''

02:40 preset failed when we were slewing back to the target after obtaining the sky frames

02:44 repeat preset

02:50 IC10 dither pattern in Ks

03:10 preset cancelled because we manually tried to offset

03:25 IC10 dither pattern in Ks

03:48 this time the AO system managed to close the loop when returning to the target after the sky

03:50 starting IC10 dither pattern in Ks

04:15 starting IC10 dither pattern in J

04:50 preset cancelled - the objects has lost 2mag - clouds?
  • going out no clouds are obvious ...

05:17 preset IC10 just to check if the cloud is gone...
  • seeing is also getting worse - dimm seeing 2.0''
  • preset cancelled - the ref star seems too faint
  • the temperature has dropped 2C in half an hour

05:22 slew to NGC1560
  • failed
  • dimm seeing >2.0''
  • in the meantime we switch on mirror ventilation becuase ambient temperature has dropped 2C

05:49 since seeing is so bad, we slew to HD15475
  • the preset keeps failing giving as error that the star is too bright

06:00 try pointing correction and collimation
  • from the guider we see that the seeing is getting better - we slew back to the original targets
  • the temperature is stabilizing

06:10 slew to NGC1560
  • mirror ventilation is off
  • it failed again - reading shared buffer master diagnostics
  • restart AOS

06:25 preset NGC1560

06:40 after restarting the AOS - still the commands are failing
  • restart the arbitrator - if the software is restarted the arbitrator has to be started separately

06:49 loop is finally closed!!!

06:51 preset cancelled - unknown reason
  • repeat preset
  • dimm seeing 1.2''

06:58 dither NGC1560 - Ks (12:00MST)
  • the loop opened because of variable seeing
  • stop AO - start AO - success
  • we did not need to interrupt the dither pattern
  • 18min in Ks

07:27 dither NGC1560 - J (12:27MST)

07:30 preset cancelled - guiding is not active!
  • repeat preset - preset complete
  • preset failed immediately after it gave success ....

07:36 preset NGC1560

07:52 dither NGC1560 - J

08:35 dither NGC1560 - K

09:05 preset cancelled
  • communication problem with ccd39
  • to recover error reboot AOS and switch on arbitrator

09:26 during preset we get a completely defocused guide star - actually we get a white orizzontal line...
  • check pointing and collimation

10:03 Astigmatism is very difficult to correct -still trying to resolve
  • they are on skype with Marco in Arcetri
  • problem identified - the astigmatism look up table has not been loaded since the last run

10:19 preset NGC1560
  • failed

10:22 checking collimation on a bright star sending the preset directly from the telescope iraf
  • the first time the preset failed - the M2 hexapod had to be initialized

10:35 Clearing all the optics - the problem seems to be on the side of the telescope
  • the star is vibrating on the guider - we cleared all the optics - we reset the probe but the star is still jumping in and out of the guider
  • the star is stable on the ccd47
  • call John Hill and A. Rakich

10:50 restarted GCS - now the star is not bouncing around.
  • problem seems solved - 2hr lost!!!!!
  • it was difficult to diagnose this problem because it had not been seen before.

11:05 preset NGC1560

11:12 repeat preset

11:23 pointing correction and collimation

11:30 preset QSO0848 - failed
  • we still see high abberations
  • flat the shell - clear optics - manually insert +2000 Z7

11:47 now the active optics seem converged
  • preset QSO0848 - failed while closing the loop with 10 modes

11:52 preset QSO0848
  • still problems with the AO secondary - it does not manage to close the loop

12:23 preset QSO0848

12:32 observe routine exited with error

12:36 Slew to BS9142
  • the WFS is not updating the rms - no corrections sent from the active optics
  • we sent preset again and the star is completely out of focus - the preset failed
  • we ask the TO to collimate on BS9142

12:55 Preset BS9142 again
  • the WFS is not sending corrections - rms =0
  • apparently the GS is mag=10 and the WFS is saturated
  • we obtained the standard star in 9 positions in Ks and J with exptime 30sec and 60sec

13:50 WFS saturated - preset cancelled because we are well within twilight

13:54 going to zenith for flats in J and Ks

14:33 Close dome - take dome flats in J and Ks

15:00 END OF NIGHT

-- KonstantinaBoutsia - 30 Nov 2011
Topic revision: r7 - 30 Nov 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback